# Modelling bulk mechanical effects in a planar cellular monolayer

### v1

@@ Appendix D Height variation across a constrained monolayer
$U={\textstyle\sum_{h}}\mathcal{U}(A_{h},L_{h},H_{h})+\tfrac{1}{2}\Lambda\mathsf{H}^{\top}\mathcal{L}\mathsf{H},$ (42)
@@ Appendix E Asymptotic limits
$L\approx\frac{1}{H},~{}~{}A\approx\frac{\sqrt{3}}{24H^{2}},~{}~{}T\approx-\tfrac{1}{2}\Gamma_{A}+\frac{\sqrt{3}}{24H},$ (46a)
$P^{2D}\approx-H,~{}~{}T^{2D}\approx\frac{\sqrt{3}}{12},\quad L_{0}\approx-\frac{\sqrt{3}}{48\Gamma_{L}},$ (46b)
with $H\gg 1$. Here the tension $T$ balances adhesive forces. Floppy states under large $\Gamma_{a}$ are obtained by solving (45) with $P^{2D}=0$, $T^{2D}=0$ for $A$, $H$, $L$ and $T$. This limit is shown on Fig. 3(a,c). Imposing (19) in addition yields, for $\Gamma_{a}\gg 1$ the limits shown in Fig. 6(b,d,e).
@@
00 0 $=A(AH-1)+(\Gamma_{a}(2A+L_{0}H-1)+\tfrac{1}{2}\Gamma_{A})L_{0},$ (47c)
@@
In this case, the hexagonal constraint (47a) determines $A$, the vertical force balance (47c) specifies $H$ and the Lagrange multiplier $T^{c}$ takes whatever value is needed to satisfy the force balance (16a, b). This leads to asymptotes shown in Fig. 5. Floppy solutions satisfying (17) in this limit are recovered by solving (47b,c,d) with $P^{2D}=0$, $T^{2D}=0$ for $A$, $H$ and $T^{c}$ (and $L=L_{0}$). Imposing (19) in addition leads to limits plotted in Fig. 6(c,d,f). Setting $H=0$ in (47a,c) leads to
$\Gamma_{A}=\frac{\sqrt{3}}{12}L_{0}+\Gamma_{a}\left(2-\frac{\sqrt{3}}{6}L_{0}^{2}\right),$ (48)
while (47a,c) in the limit $H\rightarrow\infty$ requires
$\Gamma_{a}=-L_{0}^{3}/192.$ (49)
@@
$A\approx\frac{96\Gamma_{a}^{2}}{\sqrt{3}\Gamma_{A}^{2}},\quad L\approx\frac{48\Gamma_{a}}{\sqrt{3}(-\Gamma_{A})},\quad H\approx\frac{\sqrt{3}\Gamma_{A}^{2}}{96\Gamma_{a}^{2}},$ (52)
@@ Appendix F Dynamics
Figure 11: Eigenmodes (a) and eigenvalues of the Hessian $\mathcal{H}$ as (b) $L_{0}$, (c) $\Gamma_{a}$ and (d) $-\Gamma_{A}$ are varied, for a hexagon at baseline parameter values (Table 1). Eigenvalues correspond to the mode(s) of the same colour in (b–d); modes of the same colour in (a) share an eigenvalue. The dilation mode is shown in cyan. Translation and rotation modes (with zero eigenvalue) are not shown.
[…]
@@ Appendix H Stiffness
These coefficients contain second derivatives of the energy, contributing to material stiffness, plus first derivatives, contributing to geometric stiffness. The first derivatives can be identified as prestresses in the base state in (59). Eq. (60) demonstrates how the cross-coupling in $\mathcal{U}$ leading to non-zero mixed second derivatives can influence the cell stiffness.
@@
=$=\big{\{}(2A^{2}\mathcal{U}_{AA}+AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A})\mathsf{I}_{\perp}$
@@
+(2+(2+ ( 2 $AH\mathcal{U}_{AH}+HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}-2H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (61a)
@@
=$=\big{\{}(2A^{2}\mathcal{U}_{AA}+2AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A}$
@@
+(2+(2+ ( 2 $AH\mathcal{U}_{AH}+HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}-2H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (62b)
@@
=$=\big{\{}(-A^{2}\mathcal{U}_{AA}-\tfrac{1}{2}AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A})\mathsf{I}_{\perp}$
@@
+(+(-+ ( - $AH\mathcal{U}_{AH}-\tfrac{1}{2}HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}+H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (65)
@@
=$=\big{\{}(-A^{2}\mathcal{U}_{AA}-AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A}$
@@
+(+(-+ ( - $AH\mathcal{U}_{AH}-\tfrac{1}{2}HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}+H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e,$ (66)

### v2

@@ Appendix D Height variation across a constrained monolayer
$U={\textstyle\sum_{h}}\mathcal{U}(A_{h},L_{h},H_{h})+\tfrac{1}{2}\Lambda\mathsf{H}^{\top}\mathcal{L}\mathsf{H}$ (42)
@@ Appendix E Asymptotic limits
$L\approx\frac{1}{H},~~A\approx\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{24H^{2}},~~T\approx-\tfrac{1}{2}\Gamma_{A}+\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{24H},$ (46a)
$P^{2D}\approx-H,~~T^{2D}\approx\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{12},\quad L_{0}\approx-\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{48\Gamma_{L}},$ (46b)
with $H\gg 1$. Here the tension $T$ balances adhesive forces. Floppy states under large $\Gamma_{a}$ are obtained by solving (45) with $P^{2D}=0$, $T^{2D}=0$ for $A$, $H$, $L$ and $T$, which accommodates the solution for elongated cells
$A=0,\quad H=1/L_{0},\quad L=L_{0},\quad T=-\tfrac{1}{2}\Gamma_{A},$ (47)
shown on Fig. 3(a,c). Imposing (19) on floppy states yields the large-$\Gamma_{a}$ limits shown in Fig. 6(b,d,e) and 14(a).
@@
$0$ $=A(AH-1)+(\Gamma_{a}(2A+L_{0}H-1)+\tfrac{1}{2}\Gamma_{A})L_{0},$ (48c)
@@
In this case, the hexagonal constraint (48a) determines $A$, the vertical force balance (48c) specifies $H$ and the Lagrange multiplier $T^{c}$ takes whatever value is needed to satisfy the force balance (16a, b). This leads to asymptotes shown in Fig. 5. Floppy solutions satisfying (17) in this limit are recovered by solving (48b,c,d) with $P^{2D}=0$, $T^{2D}=0$ for $A$, $H$ and $T^{c}$ (and $L=L_{0}$). Imposing (19) in addition leads to limits plotted in Fig. 6(c,d,f) and 14(c). Setting $H=0$ in (48a,c) leads to
$\Gamma_{A}=\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{12}L_{0}+\Gamma_{a}\left(2-\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}}{6}L_{0}^{2}\right),$ (49)
while (48a,c) in the limit $H\rightarrow\infty$ requires
$\Gamma_{a}=-L_{0}^{3}{\color[rgb]{0,0,0}a_{0}^{6}}/192.$ (50)
@@
$A\approx\frac{96\Gamma_{a}^{2}}{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}\Gamma_{A}^{2}},\quad L\approx\frac{48\Gamma_{a}}{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}(-\Gamma_{A})},\quad H\approx\frac{\sqrt{3}{\color[rgb]{0,0,0}a_{0}^{3}}\Gamma_{A}^{2}}{96\Gamma_{a}^{2}},$ (53)
@@
Constrained volume. Here we consider
$H$ $=1/A,$ (54a)
$A$ $=\sqrt{3}a_{0}^{3}L^{2}/24,$ (54b)
$0$ $=AP+[\Gamma_{a}(2A+LH-1)+\tfrac{1}{2}\Gamma_{A}]L,$ (54c)
$P^{2D}$ $=HP+2\Gamma_{a}(2A+LH-1),$ (54d)
$T^{2D}$ $=\Gamma_{a}(2A+LH-1)H+\tfrac{1}{2}\Gamma_{A}H+2\Gamma_{L}(L-L_{0}),$ (54e)
$0$ $=2AP^{2D}+LT^{2D}$ (54f)
in the limit $\Gamma_{a}\sim\Gamma_{A}\sim\Gamma_{L}\ll 1$. We include $P$ as a Lagrange multiplier enforcing the volume constraint $AH=1$. (Floppy states satisfy (54a,c,d,e) with $P^{2D}=T^{2D}=0$ and (18).) Setting $\Gamma_{L}=0$, scaling $P$, $P^{2D}$ and $T^{2D}$ on $\Gamma_{A}$ and setting $\hat{\Gamma}_{A}=\Gamma_{A}/\Gamma_{a}$, (54) reduces to
$(9l^{3}-2)(\sqrt{3}(9l^{3}+4)-3a_{0}l)=3\hat{\Gamma}_{A}la_{0}$ (55)
where $l=\frac{1}{6}a_{0}^{2}L$. The left-hand-side of (55) is a product of two cubics: the first has a single zero at $\ell=(2/9)^{1/3}$ in $\ell>0$; the second cubic (proportional to 2A+LH12A+LH-1) has two positive zeros for $a_{0}>2^{2/3}3^{7/6}\approx 5.71$. These zeros also first appear at $\ell=(2/9)^{1/3}$, implying that $\ell$ becomes a multivalued function of $\gamma$ for $a_{0}>2^{2/3}3^{7/6}$. Floppy states for $\Gamma_{L}=0$ and $a_{0}>2^{2/3}3^{7/6}$ exist along $\hat{\Gamma}_{A}=0$, connecting roots of (55) for which $2A+LH=1$. Thus, as predicted by [31] using a slightly different model, the squamous-to-columnar transition with $\Gamma_{L}=0$ is abrupt and takes place at $\hat{\Gamma}_{A}=0$.
Figure 11: Example illustrating a sharp squamous-to-columnar transition when cell volume is constrained. Green curves show solutions of (54) for $a_{0}=7$, $\hat{\Gamma}_{L}\equiv\Gamma_{L}/\Gamma_{a}=0.1$ for varying $\hat{\Gamma}_{A}\equiv-\Gamma_{A}/\Gamma_{a}$. Orange lines show floppy states satisfying (54a,c,d,e) with $P^{2D}=T^{2D}=0$ for which $A\leq\sqrt{3}a_{0}^{3}L^{2}/24$. Symbols indicate the two bifurcation points where the floppy and rigid solution branches intersect and rigid states lose stability (solid curves become dashed).
For non-zero $\Gamma_{L}$, however, the branch of floppy solutions bends away from $\Gamma_{A}=0$ and intersects the branches of rigid solutions at new locations. Fig. 11 illustrates one example. The transition is continuous but a branch of floppy states connects the locally stable rigid squamous and columnar states.
@@ Appendix F Dynamics
Figure 12: Eigenmodes (a) and eigenvalues of the Hessian $\mathcal{H}$ as (b) $L_{0}$, (c) $\Gamma_{a}$ and (d) $\Gamma_{A}$ are varied, for a hexagon at baseline parameter values (Table 1). Eigenvalues correspond to the mode(s) of the same colour in (b–d); modes of the same colour in (a) share an eigenvalue. The dilation mode is shown in cyan. Translation and rotation modes (with zero eigenvalue) are not shown.
@@
Figure 12 shows the values of the non-zero eigenvalues of $\mathcal{H}$ for an isolated hexagonal cell, for varying $L_{0},\Gamma_{a}$ and $\Gamma_{A}$, as discussed in Sec. III. In all cases, all symmetry-breaking modes simultaneously become unstable when $T^{2D}=0$. The cell height vanishes at $\Gamma_{A}\approx 1.42$ when $\Gamma_{A}$ is varied. All other unstable points correspond to a loss of in-plane rigidity. Eigenvalues of the symmetry-breaking modes diverge to $\pm\infty$ when the equilibrium cell area and perimeter go to zero.
@@ Appendix H Stiffness
These coefficients contain second derivatives of the energy, contributing to material stiffness, plus first derivatives, contributing to geometric stiffness. The first derivatives can be identified as prestresses in the baseline state in (62). Eq. (63) demonstrates how the cross-coupling in $\mathcal{U}$ leading to non-zero mixed second derivatives can influence the cell stiffness.
@@
$=\big{\{}(2A^{2}\mathcal{U}_{AA}+AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A})\mathsf{I}_{\perp}$
@@
$+(2$ $AH\mathcal{U}_{AH}+HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}-2H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (64a)
@@
$=\big{\{}(2A^{2}\mathcal{U}_{AA}+2AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A}$
@@
$+(2$ $AH\mathcal{U}_{AH}+HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}-2H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (65b)
@@
$=\big{\{}(-A^{2}\mathcal{U}_{AA}-\tfrac{1}{2}AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A})\mathsf{I}_{\perp}$
@@
$+($- $AH\mathcal{U}_{AH}-\tfrac{1}{2}HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}+H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e.$ (68)
@@
$=\big{\{}(-A^{2}\mathcal{U}_{AA}-AL\mathcal{U}_{AL}+AH\mathcal{U}_{AH}-A\mathcal{U}_{A}$
@@
$+($- $AH\mathcal{U}_{AH}-\tfrac{1}{2}HL\mathcal{U}_{HL}+H^{2}\mathcal{U}_{HH}+H\mathcal{U}_{H})\hat{\mathbf{z}}\otimes\hat{\mathbf{z}}\big{\}}e,$ (69)
@@
Figure 13: Impact of varying $a_{0}$, presented following the format of Fig. 2, except that geometric variables are plotted without implementing the rescaling in (13).
Appendix I Target surface area
The effects of varying the parameter $a_{0}$ are shown in Figs 13 and 14. Both figures show variables and parameters without implementing the scaling by $a_{0}$ in (13). Axes and legend labels are given in terms of the scaled values used elsewhere but with the tilde explicitly shown here.
Figure 14: Cross-sections of $a_{0}$ with $(L_{0},\Gamma_{a},-\Gamma_{A},\Gamma_{L})$-parameter space, intersecting the baseline case (black circle). Parameters are plotted without implementing the rescaling in (13). In each plot, only two parameters are varied. Hatching shows regions where there are no physical solutions. Grey regions show the floppy regime. Blue solid lines show where $T^{2D}=0$, corresponding to a rigidity transition. Blue dashed lines show asymptotes (45) and (19) with $P^{2D}=T^{2D}=0$ in (a), and (48) with $P^{2D}=T^{2D}=0$ in (c). The black dashed line in (a) shows (47).